OpenStudy (anonymous):

Sometimes you'll need to simplify more complicated logarithmic equations before using the strategy you learned. For example, to solve an equation of the form a*logb x = d, you should first divide both sides of the equation by the coefficient _____.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

the coefficent of a term is the constant that it is being multiplied by. so in this case, the term on the left is being multiplied by a, so this is your coefficent.

OpenStudy (anonymous):

alogb x = d (assuming you mean log x base b): logb x = d/a logx/logb = d/a logx = (d/a) * logb x = 10^((d/a) * logb)